Two cars leave towns 230 km apart at the same time, traveling directly toward one another. One car travels 15km per hr slower than the other. They pass one another 2 hours later. What are their speeds?
Answer by checkley71(8403)

230=2(X)+2(X-15) OR 230=2X+2X-30 OR 4X=265 OR X=50KM
PROOF 230=2*65+2*50 OR 230=130+100 OR 230=230